OpenStudy (anonymous):

Which of the following is the solution to the equation 8(y + 2) = 2y ? y = 1 y = 3 y = –3 y = –1

OpenStudy (lgbasallote):

distribute 8 into the parenthesis 8y + 16 = 2y now put all y on one side 8y - 2y + 16 = 0 put 16 on the other side 8y - 2y = -16

OpenStudy (lgbasallote):

are those options right?? o.O

OpenStudy (anonymous):

yes